A golf ball is driven with a velocity of 60 m/s at an angle of 37° above the horizontal. It strikes a green at a horizontal distance of 250m from the tee. What was the elevation of the green above the tee?


1
Expert's answer
2021-06-21T05:52:39-0400

elevation of the green above the tee "\frac{V_0^2(sin\theta)^2}{2g} = \frac{60^2\times sin30^2}{2\times9.81}=45.871m"